OpenStudy (anonymous):

2-v over -2 = 3v-2 over 5 ((help solve))

OpenStudy (gorv):

\[2-\frac{ v }{ -2}=3v- \frac{ 2 }{ 5 }\]

OpenStudy (gorv):

-and - will become plus \[2+\frac{ v }{ 2 }=3v-\frac{ 2 }{ 5 }\]

OpenStudy (gorv):

subtract v/2 om both side

OpenStudy (anonymous):

then ?

OpenStudy (gorv):

what u got ??

OpenStudy (gorv):

add 2/5 on both side

OpenStudy (anonymous):

wait where do i subtract v/2?

OpenStudy (gorv):

\[2+\frac{ v }{ 2 }=3v-\frac{ 2 }{ 5 }\] from both side

OpenStudy (anonymous):

so from v/2 and 2/5 ?

OpenStudy (gorv):

subtract v/2 from both side first

OpenStudy (anonymous):

so i subtract it from 2 and 3v?

OpenStudy (gorv):

\[2+\frac{ v}{ 2}-\frac{ v }{ 2 }=3v-\frac{ 2 }{ 5 }-\frac{ v }{ 2 }\]

OpenStudy (gorv):

what u got after solvin??

OpenStudy (anonymous):

2 = 3v - 2(v-5) over 5v ?

OpenStudy (wolf1728):

I thought the problem would be set up like this: